Physics help!!
Q 4
An object is placed 40 cm in front of a concave wi
Answer #1

Option D) is correct.

N(t) = No*exp[-decay constant*time] = 1.85*10^12*exp[-1.6*10^-3*30*60] = 1.04*10^11